OpenStudy (anonymous):

re-write this equation in standard form : y-3=2.5(x+1)

OpenStudy (anonymous):

The standard form is Ax + By + C = 0 multiply all terms by 2 2y -6 = 5(x + 1) multiply out the bracket 2y - 6 = 5x + 5 and rearrange 5x - 2y +11 = 0